Hitachi – The Reliability Leader in Spectroscopy
Who has a Beckman 600 series filter wheel?
I have a Beckman 650 and the filter wheel appears to have one "bad" filter on it (a near UV that has frosting or pitting on one side of the glass). I suspect it may be a standard filter glass. The machine still works, but this lowers the UV energy so that it will no longer run the wavelength cal routine. (note; this is normally run on startup and at the same time it automatically runs SETCALGAIN. Since my machine was not running the first, the second never ran and I had problems with doing any scans. Running SETCALGAIN in diagnostics fixed that).
Does anybody have a dead 600 series machine from which I can get or buy the filter wheel? I can then measure the transmission curve of this filter and find out if it is a std filter glass for replacement. I suspect I am not the only one to have this glass go bad, so I can make a cut to size filter available for others to fix their machines (assuming it all works as I hope!).
Or does anybody have a list of filter types used in the wheel?
Before changing the filter wheel or filter, have you cleaned it? Are you getting errors during startup?
I have a filter wheel I can sell you. But, I've worked on these for over 20 years and I haven't ran across a filter problem. Not saying it couldn't happen though.
From a cold startup it passes all the power up diagnostics.
After starting and warming up both lamps and starting diagnostics to run "calibrate wavelength" I get a Wavelength Cal error: Zero order.
One filter on the filter wheel has a frosted appearance on one side (the others are transparent); is this normal? I tried cleaning this filter only with alcohol; when wet it looks close to transparent, but returns to frosted after drying. No improvement was noted. It is a bluish/dark filter.
Zero Order error is saying it can't find white light when it goes to 0nm, this is the first thing it checks when you do a calibrate wavelength.
In readiread set wavelength to 0nm. Put a peice of white paper in front of the detector (right side of the sample compartment), you should see a white square. It should be fairly bright. At 0nm the light path should be going through the slit (before the filter wheel) and through a blank spot (or clear filter) in the filter wheel. If the light is dim, make sure the lamp is peaked. Follow the light path with a peice of white paper and see where the light is getting dim. I would recommend DO NOT clean the grating mirror. It looks foggy, but that is the way the mirror is designed. It can be cleaned, but be very careful not to scratch it or leave any residue.
Hope this helps, good luck.
Thanks for the extremely valuable advice! I found that when sitting at zero nm, the bright stripe of light from the diffraction grating mirror was about 2mm to the right of the metal slit. Since the slit was only taped into place (looks like glued had failed sometime in the past), I un-taped, moved it onto the beam and the light went through!
I now did a setscangain, then wavelength cal..... and it actually completed. But when I put the cover back on wavelength cal failed. Pulling the cover off again and trying several more times both on and off I got fails. I did see that during wavelength cal it uses a couple filters initially, then goes to each filter. I suspect that is where my "frosted " filter is causing me trouble. Viewing from the motor side of the filter wheel, and counting from the zero lug. The second filter (clockwise) is frosted. I measured the transmission and it appears to match a Hoya B390.
I had the filter wheel out and did a transmission scan (using my trusty older BeckmanDU64!) of the frosted blue filter (it is in the #2 position, going clockwise, when looking at it from the motor side) and it seems to be a good match to a Hoya B-390 (I got 390 or 395 center peak and 112nm BW, the Hoya B390 is 390nm center and 125 nm BW). The filter in the wheel seems to be ~2.4mm thick so it looks like Edmund has 2.5mm thick Hoya B390 for sale as std filter material.
As a test, I have applied a very small amount of Nye Delicate machine oil #113 to the frosted surface of my problem filter (to make it much more transparent). Darn... it still fails wavelength cal. I cleaned the oil off the filter (oddly enough it did not seem to come off , even with acetone, lens cleaner, and alcohol and now it is much clearer than it used to be. One benefit of the filter being clearer is that I have much reduced scan noise in the 357 to 412 nm range.
After all that I now notice that the original problem; nm set at Zero, should see nice bright light is bad because the aperture is not lined up. Maybe due to drift of warmup? I don't think the tape moved that much (about a mm), but it is possible. Tried SETSCANGAIN and Wavelength Cal again ; failed. Noticed that the height of the slit was off (should set it with both zero nm and at 550nm). Moved slit to centered in both axis; tried SETSCANGAIN and Wavelength cal again;fail,
I have noticed that the beam is kissing the back edge of the sensor hole/window, I rotated the nearest mirror so that the beam is centered in the sensor. Still no good wavelength cal.
I finally looked in service mode and see a very nice table of scan gains. All from 190nm to 415nm are listed as 255. from 415.1 to 1100nm are 80, 40, 56. Seems like this might be a UV bulb issue. I will next clean all the filters and mirrors (NOT the grating!) since UV energy dissipates fast with films and dirt.
My wavelength now calibrates! It seems that someone had moved the exit slit, so the lamp/slit alignment was loosing me a lot of light (my UV lamp ia also probably very weak). Here is what I did;
To align lamps;
1) Use RediRead to set at 200nm
2) Align UV by loosening 4 screws and sliding the casting holding both lamps left-right for peak reading. If you have the optical cover off, center the violet circle on the entrance slit. Tighten 4 screws.
3) Use RediRead to set at 500nm
4) Align white lamp-loosen 2 lamp screws and adjust lamp front-back for max reading. If you have the optical cover off, center the bright slash of light on the entrance slit.
1) With the optical cover off, use RediRead to set to 0nm. This will show a bright stripe of light on the exit slit, which should be centered on the slit. If not centered (or at least fully covering the slit), loosen the tape holding the slit and move it so that the slit is centered and parallel to the stripe of light.
2) Change RediRead to 500nm. Verify that the slit is positioned vertically so the color band is fully getting through it.
Then go to diagnostics and run SETSCANGAIN.